OpenStudy (anonymous):

the first three terms of an arithmetic are p, 2p-6 and 5p-12. what is the 4th term?

OpenStudy (anonymous):

if it is arithmetic, then the differences need to be the samdoesn't look arithmetic to me, but maybe i am missing something. e \(2p-6-p=p-6\) and \(5p-12-(2p-6)=3p-6\) so if \[p-6=3p-6\] we best conclude that \(p=0\)

OpenStudy (anonymous):

wow that came out garbled, didn't it. what i meant was if it is arithmetic then is seems that p must be 0